# algorithmic

### v3

@@ 5 Simulation results
We evaluate the performance of the proposed EKF-ADMM algorithm on three different examples: online LASSO [18], online training of a neural network on data from a static model under $\ell_{1}$ regularization or bound constraints, online adaptation of a neural network on data from a time-varying model under $\ell_{0}$ regularization.
@@ 5.1 Online LASSO
Consider the LASSO problem $\min_{x}\sum_{k=0}^{N-1}(\frac{1}{2}\left\|y_{k}^{m}\right.$- $\left.C_{k}x\right\|^{2}_{R^{-1}}$ $+$ $\lambda\left\lVert x\right\rVert_{1})$, where $x\in\mathbb{R}^{3}$ is the parameter vector, $C_{k}\in\mathbb{R}^{2\times 3}$ are randomly generated matrices with $C_{k,ij}\sim\mathcal{N}(0,1)$, $y_{k}^{m}=C_{k}x_{\rm true}+r_{k}\in\mathbb{R}^{2}$ is the vector of measurements and $r_{k}$ is random measurement noise, $r_{k,i}\sim{\mathcal{N}}(0,10^{-2})$. We will evaluate the behavior of the regret functions $R_{f}(N)$ and $R_{c}(N)$ as $N\rightarrow\infty$ when using Algorithm 2. The following settings are used: $P_{0|-1}=I,Q_{k}=10^{-6}I,R=10^{-3}I,k_{n}=10^{3},\rho=10^{4}\sqrt{N}$ and $\eta=10^{-6}\sqrt{N}$, where $P_{0|-1},Q_{k}$ and $R$ are manually tuned to maximize the Kalman filter performance without regularization and $\rho,\eta$ are chosen according to the expressions provided in Theorem 4.1. Results for different values of $\lambda$ are shown in Figure 1. In this case, Theorem 2 holds and, as expected, both the regrets $R_{f}(N)$ and $R_{c}(N)$ decrease as the number $N$ of samples increases.
@@ 5.2 Online learning of a static model
Consider the dataset generated by the static nonlinear model $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{10}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$. We want to train online a neural network $h_{k}(x)$ with $2$ layers, $8$ neurons in each layer, and $\tanh$ activation function, with $n_{x}=105$ trainable weights in total. The training is performed on $N=10^{5}$ randomly generated data points, where $z_{k,i}\sim\mathcal{U}_{[-10,10]}$ and $r_{k}\sim{\mathcal{N}}(0,10^{-2})$. Let $\{x_{k}\}_{k=0}^{N-1}$ be the sequence of weights generated by Algorithm 1: we evaluate the online adaptation performance by means of the regret function $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(x_{k}))-\min_{x}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$, where $f_{k}(x)=\frac{1}{2}\left\lVert y_{k}^{m}-h_{k}(x)\right\rVert^{2}$, and the quality of a given solution $x$ using the performance indices $\text{Loss}(x)=\frac{1}{N}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$, $\text{Mse}(x)=\frac{1}{N}\sum_{k=0}^{N-1}f_{k}$, $\text{Reg}(x)=\frac{1}{N}\sum_{k=0}^{N-1}g(x)$ and $\text{Cv}(x)=\left\lVert x-\Pi_{\mathcal{C}}(x)\right\rVert_{2}^{2}$, where, given a constraint set $\mathcal{C}\subseteq\mathbb{R}^{n_{x}}$, $\Pi_{\mathcal{C}}(x)$ is the projection of the point $x$ onto $\mathcal{C}$. The training is performed in MATLAB R2022a on an Intel Core i7 12700H CPU with 16 GB of RAM, using the library CasADi [19] to compute the required Jacobian matrices via automatic differentiation. All results are averaged over 20 runs starting from different initial conditions, that were randomly generated using the well known Xavier’s initialization strategy.
@@ 5.2.1 $\ell_{1}$ regularization
We train the neural model under the regularization function $g(x)=\lambda\left\lVert x\right\rVert_{1}$, with $\lambda=10^{-4}$. We selected the following hyper-parameters: $\rho=10\lambda$, $n_{a}=1$, $Q_{k}=10^{-4}I$, $R_{k}=I$ and $P_{0|0}=100I$. We compare the results to different online optimization alternatives: online ADMM [13] with constant matrix $P=10^{-2}I$ (online-ADMM), EKF-ADMM with time-varying $\rho_{k}=10^{\frac{k}{N}-2}\lambda$ (EKF-ADMMtv), EKF with $\ell_{1}$-regularization [8] (EKF-$\ell_{1}$) and SMIDAS [20] with learning rate $\eta=5\cdot 10^{-2}$. Notice that $P_{0|0},R_{k}$ and $Q_{k}$ are chosen by manually tuning EKF-$\ell_{1}$ and then used for all other approaches, while all the remaining method-specific hyperparameters have been manually tuned to maximize their performance. The reason for choosing a time-varying $\rho_{k}$ is that fake measurements are usually not accurate initially, so that it is better to start with a higher value of $\frac{1}{\rho_{k}}$ and then decrease it progressively. In addition, we compare with two offline batch algorithms: NAILM [21] and LBFGS [22], the latter using the Python library jax-sysid. Note that all the online algorithms consume the dataset only once (1 epoch), except NAILM and LBFGS that run over 150 and 5000 epochs respectively. The results obtained at the end of the training phase are reported in Table 1.
@@
online-ADMM [13] $10.38$ ($1.7$) $4.68$ ($1.8$) $3.62$ ($2.21$) $530.69$ ($29.09$)
EKF-$\ell_{1}$ [8] $5.47$ ($0.67$) $1.42$ ($0.26$) $56.42$ ($7.67$) $12.46$ ($0.27$)
SMIDAS [20] $89.93$ ($36.3$) $84.67$ ($36.6$) $61.71$ ($9.98$) $4.11$ ($1.11$)
@@
Figure 2: Online learning with $\ell_{1}$ regularization: Loss, Mse and Reg averaged over $20$ runs.
Figure 3: Online learning with $\ell_{1}$ regularization: regret and sample regret averaged over $20$ runs.
@@ 5.2.2 Bound constraints
Let us now repeat the training under the bound constraints imposed by the regularization function $g(x)=0$ if $x\in\mathcal{C}$ and $g(x)=+\infty$ otherwise, where $\mathcal{C}=\{x\in\mathbb{R}^{n_{x}}:\ |x_{i}|\leq 0.5\}$. We use the hyper-parameters $\rho=1,n_{a}=5,Q_{k}=10^{-4}I,P_{0|0}=100I$, and $R_{k}=I$. In this example, we also compare with a simple clipping step of the Kalman filter (EKF-CLIP) and an online projected gradient method (OGD-proj) [4] with learning rate $\eta=10^{-3}$. Matrices $P_{0|0},R_{k}$ and $Q_{k}$ are the same as for the $\ell_{1}$ example and are in common for all the Kalman-like approaches, while all the other hyperparameters have been manually re-tuned to optimize performance in the bounds constraints example. Results obtained at the end of the training phase are reported in Table 2. Among the online approaches, considering the final Mse, Cv and execution time, EKF-ADMM provides the best quality solution, while OGD-proj attains similar performance but at a slower pace. Figure 4 shows the performance of the solution during the training phase.
We have noticed that the choice of $\rho$ is crucial to obtain good adaptation performance: while for some values the convergence is slow and not satisfactory, with limited effort and a proper manual tuning, we were easily able to find values with good performance in all the examples that we have shown here.
@@
online-ADMM [13] $0.129$ ($0.010$) $90.77$ ($59.37$) $610.73$ ($9.82$)
EKF-CLIP $0.214$ ($0.048$) 00 (00) $11.89$ ($0.12$)
OGD-proj [4] $0.137$ ($0.013$) 00 (00) $2.56$ ($0.15$)
@@ 5.3 Online learning of a time-varying model
We test now the ability of EKF-ADMM to adapt the same neural network model, under $\ell_{0}$ regularization, when the data-generating system switches as follows: $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{10}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $k\leq\frac{N}{3}$, $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{2}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $\frac{N}{3}<k\leq\frac{2N}{3}$ and $y^{m}_{k}=\frac{0.3\cdot z_{k,1}^{2}-e^{\frac{z_{k,2}}{2}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $\frac{2N}{3}<k$, with $N=1.5\cdot 10^{5}$, $z_{k,i}\sim\mathcal{U}_{[-10,10]}$ and $r_{k}\sim{\mathcal{N}}(0,10^{-2})$. We evaluate the regret function $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(x_{k}))-\min_{z_{1},z_{2},z_{3}}\sum_{i=1}^{3}r_{i}(z_{i})$, with $r_{i}(z_{i})=\sum_{k=(i-1)\frac{N}{3}}^{i\frac{N}{3}}(f_{k}(z_{i})+g(z_{i}))$, where $\{x_{k}\}_{k=0}^{N-1}$ is the sequence generated by Algorithm 1. The regularization term is $g(x)=\lambda\left\lVert w\right\rVert_{0}$, with $\lambda=10^{-4}$, and we use the EKF-ADMM hyper-parameters $\rho=10^{3}\cdot\lambda$, $p_{a}=1$, $Q_{k}=10^{-4}I$, and $P_{0|0}=100I$, which were manually tuned by optimizing the EKF performance. Since the model is now time-varying, we will also use an EKF implementation with forgetting factor $\alpha=0.9$ [9]. The resulting regret function is shown in Figure 5. It is apparent that EKF-ADMM can effectively track changes of the underlying data-generating system.
Figure 5: Online learning with $\ell_{0}$ regularization: regret and sample regret averaged over $20$ runs.
@@ 6 Conclusions
Acknowledgments
This work has received support from the European Research Council (ERC), Advanced Research Grant COMPACT (Grant Agreement No. 101141351). The research work of Lapo Frascati has been financially supported by ODYS S.r.l.
@@ Appendix A Proof of Theorem 1
Starting from $R_{f}(N)$, since $x_{k+1},\nu_{k+1}$ are the optimal solutions of the first two optimization problems in Algorithm 2 and since $w_{k}=w_{k+1}-\rho(x_{k+1}-\nu_{k+1})$ we have that $\nabla f_{k}(x_{k+1})=-(w_{k+1}-\rho(\nu_{k}-\nu_{k+1}))-\eta(P_{k}^{-1}x_{k+1}-P_{k}^{-1}x_{k})$ and $w_{k+1}\in\partial g(\nu_{k+1})$, where $\partial g$ is the subgradient of $g$. Due to the convexity of $f_{k}(\cdot)$ and $g(\cdot)$, $f_{k}(x_{k+1})-f_{k}(x^{\star})\leq\nabla f_{k}(x_{k+1})^{T}(x_{k+1}-x^{\star})=-w_{k+1}^{T}(x_{k+1}-\nu^{\star})+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2}+\left\lVert x_{k+1}-\nu_{k+1}\right\rVert^{2}-\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2})+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}-$ $\left\lVert x^{\star}-x_{k+1}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{P_{k}^{-1}})$, where the equality derives from a succession of square completions and by recalling that $x^{\star}=\nu^{\star}$, and $g(\nu_{k+1})-g(\nu^{\star})\leq w_{k+1}^{T}(\nu_{k+1}-\nu^{\star})$. Summing the two inequalities together and noticing that $-w_{k+1}^{T}(x_{k+1}-\nu_{k+1})+\frac{\rho}{2}\lVert x_{k+1}-\nu_{k+1}\rVert^{2}=\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})$, we obtain:
$f_{k}(x_{k+1})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star}))\leq\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})-\frac{\rho}{2}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}+$ (14)
$+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x^{\star}-x_{k+1}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{P_{k}^{-1}})$
@@
Considering that $f_{k}(x_{k})-f_{k}(x_{k+1})\leq\nabla f_{k}(x_{k})^{T}(x_{k}-x_{k+1})\leq\frac{1}{2\alpha\eta}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\alpha\eta}{2}\left\lVert x_{k}-x_{k+1}\right\rVert^{2}$, where the second inequality is due to Fenchel-Young’s inequality, and considering Assumption A1.a of the theorem, we have that $f_{k}(x_{k})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star}))\leq\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})-\frac{\rho}{2}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})+\frac{1}{2\alpha\eta}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x^{\star}-x_{k+1}\right\rVert_{P_{k}^{-1}}^{2})$. Summing from 0 to $N-1$ and considering Assumption A3 we get
$R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star})))+g(\nu_{0})-g(\nu_{N})\leq$
@@
$\leq\frac{1}{2\rho}(\left\lVert w_{0}\right\rVert^{2}-\left\lVert w_{N}\right\rVert^{2})+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{0}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{N}\right\rVert^{2})+\frac{1}{2\alpha\eta}\sum_{k=0}^{N-1}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}\left\lVert x^{\star}-x_{0}\right\rVert^{2}_{P_{0}^{-1}}+$

### v4

@@ 5 Simulation results
We evaluate the performance of the proposed EKF-ADMM algorithm on three different examples: online LASSO [bib22], online training of a neural network on data from a static model under $\ell_{1}$ regularization or bound constraints, online adaptation of a neural network on data from a time-varying model under $\ell_{0}$ regularization.
@@ 5.1 Online LASSO
Consider the LASSO problem $\min_{x}\sum_{k=0}^{N-1}(\frac{1}{2}\left\|y_{k}^{m}\right$. - $\left.C_{k}x\right\|^{2}_{R^{-1}}$ ++ $\lambda\left\lVert x\right\rVert_{1})$, where $x\in\mathbb{R}^{3}$ is the parameter vector, $C_{k}\in\mathbb{R}^{2\times 3}$ are randomly generated matrices with $C_{k,ij}\sim\mathcal{N}(0,1)$, $y_{k}^{m}=C_{k}x_{\rm true}+r_{k}\in\mathbb{R}^{2}$ is the vector of measurements and $r_{k}$ is random measurement noise, $r_{k,i}\sim{\mathcal{N}}(0,10^{-2})$. We will evaluate the behavior of the regret functions $R_{f}(N)$ and $R_{c}(N)$ as $N\rightarrow\infty$ when using Algorithm 2. The following settings are used: $P_{0|-1}=I,Q_{k}=10^{-6}I,R=10^{-3}I,k_{n}=10^{3},\rho=10^{4}\sqrt{N}$ and $\eta=10^{-6}\sqrt{N}$, where $P_{0|-1},Q_{k}$ and $R$ are manually tuned to maximize the Kalman filter performance without regularization and $\rho,\eta$ are chosen according to the expressions provided in Theorem 4.1. Results for different values of $\lambda$ are shown in Figure 1. In this case, Theorem 2 holds and, as expected, both the regrets $R_{f}(N)$ and $R_{c}(N)$ decrease as the number NN of samples increases.
@@ 5.2 Online learning of a static model
Consider the dataset generated by the static nonlinear model $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{10}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$. We want to train online a neural network $h_{k}(x)$ with 22 layers, 88 neurons in each layer, and $\tanh$ activation function, with $n_{x}=105$ trainable weights in total. The training is performed on $N=10^{5}$ randomly generated data points, where $z_{k,i}\sim\mathcal{U}_{[-10,10]}$ and $r_{k}\sim{\mathcal{N}}(0,10^{-2})$. Let $\{x_{k}\}_{k=0}^{N-1}$ be the sequence of weights generated by Algorithm 1: we evaluate the online adaptation performance by means of the regret function $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(x_{k}))-\min_{x}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$, where $f_{k}(x)=\frac{1}{2}\left\lVert y_{k}^{m}-h_{k}(x)\right\rVert^{2}$, and the quality of a given solution $x$ using the performance indices $\text{Loss}(x)=\frac{1}{N}\sum_{k=0}^{N-1}(f_{k}(x)+g(x))$, $\text{Mse}(x)=\frac{1}{N}\sum_{k=0}^{N-1}f_{k}$, $\text{Reg}(x)=\frac{1}{N}\sum_{k=0}^{N-1}g(x)$ and $\text{Cv}(x)=\left\lVert x-\Pi_{\mathcal{C}}(x)\right\rVert_{2}^{2}$, where, given a constraint set $\mathcal{C}\subseteq\mathbb{R}^{n_{x}}$, $\Pi_{\mathcal{C}}(x)$ is the projection of the point $x$ onto $\mathcal{C}$. The training is performed in MATLAB R2022a on an Intel Core i7 12700H CPU with 16 GB of RAM, using the library CasADi [AGHRD19] to compute the required Jacobian matrices via automatic differentiation. All results are averaged over 20 runs starting from different initial conditions, that were randomly generated using the well known Xavier’s initialization strategy.
@@ 5.2.1 $\ell_{1}$ regularization
We train the neural model under the regularization function $g(x)=\lambda\left\lVert x\right\rVert_{1}$, with $\lambda=10^{-4}$. We selected the following hyper-parameters: $\rho=10\lambda$, $n_{a}=1$, $Q_{k}=10^{-4}I$, $R_{k}=I$ and $P_{0|0}=100I$. We compare the results to different online optimization alternatives: online ADMM [bib2] with constant matrix $P=10^{-2}I$ (online-ADMM), EKF-ADMM with time-varying $\rho_{k}=10^{\frac{k}{N}-2}\lambda$ (EKF-ADMMtv), EKF with $\ell_{1}$-regularization [bib1] (EKF-$\ell_{1}$) and SMIDAS [bib24] with learning rate $\eta=5\cdot 10^{-2}$. Notice that $P_{0|0},R_{k}$ and $Q_{k}$ are chosen by manually tuning EKF-$\ell_{1}$ and then used for all other approaches, while all the remaining method-specific hyperparameters have been manually tuned to maximize their performance. The reason for choosing a time-varying $\rho_{k}$ is that fake measurements are usually not accurate initially, so that it is better to start with a higher value of $\frac{1}{\rho_{k}}$ and then decrease it progressively. In addition, we compare with two offline batch algorithms: NAILM [bib4] and LBFGS [bib23], the latter using the Python library jax-sysid. Note that all the online algorithms consume the dataset only once (1 epoch), except NAILM and LBFGS that run over 150 and 5000 epochs respectively. The results obtained at the end of the training phase are reported in Table 1.
@@
online-ADMM [bib2] $10.38$ ($1.7$) $4.68$ ($1.8$) $3.62$ ($2.21$) $530.69$ ($29.09$)
EKF-$\ell_{1}$ [bib1] $5.47$ ($0.67$) $1.42$ ($0.26$) $56.42$ ($7.67$) $12.46$ ($0.27$)
SMIDAS [bib24] $89.93$ ($36.3$) $84.67$ ($36.6$) $61.71$ ($9.98$) $4.11$ ($1.11$)
@@
Figure 2: Online learning with $\ell_{1}$ regularization: Loss, Mse and Reg averaged over 2020 runs.
Figure 3: Online learning with $\ell_{1}$ regularization: regret and sample regret averaged over 2020 runs.
@@ 5.2.2 Bound constraints
Let us now repeat the training under the bound constraints imposed by the regularization function $g(x)=0$ if $x\in\mathcal{C}$ and $g(x)=+\infty$ otherwise, where $\mathcal{C}=\{x\in\mathbb{R}^{n_{x}}:\ |x_{i}|\leq 0.5\}$. We use the hyper-parameters $\rho=1,n_{a}=5,Q_{k}=10^{-4}I,P_{0|0}=100I$, and $R_{k}=I$. In this example, we also compare with a simple clipping step of the Kalman filter (EKF-CLIP) and an online projected gradient method (OGD-proj) [bib7] with learning rate $\eta=10^{-3}$. Matrices $P_{0|0},R_{k}$ and $Q_{k}$ are the same as for the $\ell_{1}$ example and are in common for all the Kalman-like approaches, while all the other hyperparameters have been manually re-tuned to optimize performance in the bounds constraints example. Results obtained at the end of the training phase are reported in Table 2. Among the online approaches, considering the final Mse, Cv and execution time, EKF-ADMM provides the best quality solution, while OGD-proj attains similar performance but at a slower pace. Figure 4 shows the performance of the solution during the training phase.
@@
online-ADMM [bib2] $0.129$ ($0.010$) $90.77$ ($59.37$) $610.73$ ($9.82$)
EKF-CLIP $0.214$ ($0.048$) 0 (0) $11.89$ ($0.12$)
OGD-proj [bib7] $0.137$ ($0.013$) 0 (0) $2.56$ ($0.15$)
@@ 5.3 Online learning of a time-varying model
We test now the ability of EKF-ADMM to adapt the same neural network model, under $\ell_{0}$ regularization, when the data-generating system switches as follows: $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{10}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $k\leq\frac{N}{3}$, $y^{m}_{k}=\frac{z_{k,1}^{2}-e^{\frac{z_{k,2}}{2}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $\frac{N}{3}<k\leq\frac{2N}{3}$ and $y^{m}_{k}=\frac{0.3\cdot z_{k,1}^{2}-e^{\frac{z_{k,2}}{2}}}{3+|z_{k,1}+z_{k,2}|}+r_{k}$ if $\frac{2N}{3}<k$, with $N=1.5\cdot 10^{5}$, $z_{k,i}\sim\mathcal{U}_{[-10,10]}$ and $r_{k}\sim{\mathcal{N}}(0,10^{-2})$. We evaluate the regret function $R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(x_{k}))-\min_{z_{1},z_{2},z_{3}}\sum_{i=1}^{3}r_{i}(z_{i})$, with $r_{i}(z_{i})=\sum_{k=(i-1)\frac{N}{3}}^{i\frac{N}{3}}(f_{k}(z_{i})+g(z_{i}))$, where $\{x_{k}\}_{k=0}^{N-1}$ is the sequence generated by Algorithm 1. The regularization term is $g(x)=\lambda\left\lVert w\right\rVert_{0}$, with $\lambda=10^{-4}$, and we use the EKF-ADMM hyper-parameters $\rho=10^{3}\cdot\lambda$, $p_{a}=1$, $Q_{k}=10^{-4}I$, and $P_{0|0}=100I$, which were manually tuned by optimizing the EKF performance. Since the model is now time-varying, we will also use an EKF implementation with forgetting factor $\alpha=0.9$ [bib3]. The resulting regret function is shown in Figure 5. It is apparent that EKF-ADMM can effectively track changes of the underlying data-generating system.
We have noticed that the choice of $\rho$ is crucial to obtain good adaptation performance. While for some values of $\rho$ the convergence is slow and not satisfactory, with limited effort, in all the reported examples, we could find appropriate values that gave good performance. We remark that, in the general nonlinear case, this might not always be the case, as both the EKF and ADMM are not guaranteed to converge, which might complicate tuning $\rho$. This could limit the use of ADMM iterations within the EKF framework in certain challenging applications.
Figure 5: Online learning with $\ell_{0}$ regularization: regret and sample regret averaged over 2020 runs.
@@ Appendix A Proof of Theorem 1
Starting from $R_{f}(N)$, since $x_{k+1},\nu_{k+1}$ are the optimal solutions of the first two optimization problems in Algorithm 2 and since $w_{k}=w_{k+1}-\rho(x_{k+1}-\nu_{k+1})$ we have that $\nabla f_{k}(x_{k+1})=-(w_{k+1}-\rho(\nu_{k}-\nu_{k+1}))-\eta(P_{k}^{-1}x_{k+1}-P_{k}^{-1}x_{k})$ and $w_{k+1}\in\partial g(\nu_{k+1})$, where $\partial g$ is the subgradient of $g$. Due to the convexity of $f_{k}(\cdot)$ and $g(\cdot)$, $f_{k}(x_{k+1})-f_{k}(x^{\star})\leq\nabla f_{k}(x_{k+1})^{T}(x_{k+1}-x^{\star})=-w_{k+1}^{T}(x_{k+1}-\nu^{\star})+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2}+\left\lVert x_{k+1}-\nu_{k+1}\right\rVert^{2}-\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2})+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}$- $\left\lVert x^{\star}-x_{k+1}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{P_{k}^{-1}})$, where the equality derives from a succession of square completions and by recalling that $x^{\star}=\nu^{\star}$, and $g(\nu_{k+1})-g(\nu^{\star})\leq w_{k+1}^{T}(\nu_{k+1}-\nu^{\star})$. Summing the two inequalities together and noticing that $w_{k+1}^{T}(x_{k+1}-\nu_{k+1})+\frac{\rho}{2}\lVert x_{k+1}-\nu_{k+1}\rVert^{2}=\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})$, we obtain:
$f_{k}(x_{k+1})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star}))\leq$ (14)
$\leq\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})-\frac{\rho}{2}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}+$
$\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert^{2}_{P_{k}^{-1}}$
$-\left\lVert x^{\star}-x_{k+1}\right\rVert^{2}_{P_{k}^{-1}}-\left\lVert x_{k+1}-x_{k}\right\rVert^{2}_{P_{k}^{-1}})$
@@
Considering that $f_{k}(x_{k})-f_{k}(x_{k+1})\leq\nabla f_{k}(x_{k})^{T}(x_{k}-x_{k+1})\leq\frac{1}{2\alpha\eta}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\alpha\eta}{2}\left\lVert x_{k}-x_{k+1}\right\rVert^{2}$, where the second inequality is due to Fenchel-Young’s inequality, and considering Assumption A1.a of the theorem, we have that $f_{k}(x_{k})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star}))\leq\frac{1}{2\rho}(\left\lVert w_{k}\right\rVert^{2}-\left\lVert w_{k+1}\right\rVert^{2})-\frac{\rho}{2}\left\lVert x_{k+1}-\nu_{k}\right\rVert^{2}+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{k}\right\rVert^{2}-\left\lVert\nu^{\star}-\nu_{k+1}\right\rVert^{2})+\frac{1}{2\alpha\eta}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}(\left\lVert x^{\star}-x_{k}\right\rVert_{P_{k}^{-1}}^{2}-\left\lVert x^{\star}-x_{k+1}\right\rVert_{P_{k}^{-1}}^{2})$. Summing from 0 to N1N-1 and considering Assumption A3 we get
$R_{f}(N)=\sum_{k=0}^{N-1}(f_{k}(x_{k})+g(\nu_{k+1})-(f_{k}(x^{\star})+g(\nu^{\star})))$
$+g(\nu_{0})-g(\nu_{N})\leq\frac{1}{2\rho}(\left\lVert w_{0}\right\rVert^{2}-\left\lVert w_{N}\right\rVert^{2})+\frac{\rho}{2}(\left\lVert\nu^{\star}-\nu_{0}\right\rVert^{2}$
@@
$-\left\lVert\nu^{\star}-\nu_{N}\right\rVert^{2})+\frac{1}{2\alpha\eta}\sum_{k=0}^{N-1}\left\lVert\nabla f_{k}(x_{k})\right\rVert^{2}+\frac{\eta}{2}\left\lVert x^{\star}-x_{0}\right\rVert^{2}_{P_{0}^{-1}}$